# Foundation Models for Time Series Analysis: A Tutorial and Survey

### v1

@@ Abstract.
${}^{{\dagger}}$. Wen is the corresponding author. Email: qingsongedu@gmail.com
@@ 2. Background
LLM, AM, VLM [General: Time-LLM (Jin et al., 2023b), OFA (Zhou et al., 2023b), LLM4TS (Chang et al., 2023), PromptCast (Xue and Salim, 2022),
TEMPO (Cao et al., 2023), LLMTime (Gruver et al., 2023), Voice2Series (Yang et al., 2021), AutoTimes (Liu et al., 2024a), leaf, text width=32em] [Finance: Yu et al. (Yu et al., 2023), Chen et al. (Chen et al., 2023f), Xie et al. (Xie et al., 2023), Wimmer et al. (Wimmer and Rekabsaz, 2023), leaf, text width=32em] [Healthcare: Liu et al. (Liu et al., 2023b), leaf, text width=22em] ] [Self-supervised [Generative [General: PatchTST (Nie et al., 2022), Moirai (Woo et al., 2024), Lag-Llama (Rasul et al., 2023),
TimeSiam (Dong et al., 2024), Timer (Liu et al., 2024c), Das et al. (Das et al., 2023), UniTS (Gao et al., 2024),
@@
[…]
Biloš et al. (Biloš et al., 2022) , Crabbé et al. (Crabbé et al., 2024) , TimeDiff (Shen and Kwok, 2023) , Wang et al.  (Wang et al., 2023e) , DiffTime (Coletta et al., 2024), leaf, text width=35em ] [Finance: FTS-Diffusion (Huang et al., 2024), leaf, text width=31.7em ] [Power: DiffLoad (Wang et al., 2023d), leaf, text width=31.7em ] ] ] [Spatial Time Series [Transformer-based [Pre-trained LLM [Transportation: ST-LLM (Liu et al., 2024b), TPLLM (Ren et al., 2024), leaf, text width=22em ] [General: GATGPT (Chen et al., 2023e), leaf, text width=22em ] ] [Self-supervised [Generative [Transportation: STEP (Shao et al., 2022), leaf, text width=22em] [Climate: W-MAE (Man et al., 2023), MetePFL (Chen et al., 2023b), FengWu (Chen et al., 2023a), leaf, text width=22em ] [General: UniST (Yuan et al., 2024), leaf, text width=22em] ] ] [Fully-supervised [Transportation: CPPBTR (Duan et al., 2019) , TFM (Wang et al., 2023c), leaf, text width=22em ] [Climate: FourCastNet (Pathak et al., 2022) , FedWing (Chen et al., 2023c) , Pangu-Weather (Bi et al., 2023a), ClimaX (Nguyen et al., 2023), leaf, text width=33em ] ] ] [non-Transformer-based
(MLP RNN CNN) [Self-supervised [General: SPGCL (Li et al., 2022b) , STGCL (Liu et al., 2022), leaf, text width=22em ] ] ] [Diffusion-based [General: DiffSTG (Wen et al., 2023a) , DSTPP (Yuan et al., 2023) , DYffusion (Cachay et al., 2023) , Yun et al. (Yun et al., 2023) , USTD (Hu et al., 2023) , PriSTI (Liu et al., 2023a), leaf, text width=42em ] ] ] [Others [Transformer-based [Pre-trained LLM [Mobility: AuxMobLCast (Xue et al., 2022) , LLM-Mob (Wang et al., 2023a), leaf, text width=22em ] [Event: LAMP (Shi et al., 2023) , Gunjal & Durrett et al. (Gunjal and Durrett, 2023), leaf, text width=22em ] ] [Self-supervised [Generative [Mobility: GTM (Lin et al., 2024), leaf, text width=22em] [Event: NYUTron (Jiang et al., 2023) , GatorTron (Yang et al., 2022), leaf, text width=22em ] ] ] ] [non-Transformer-based
@@ 4.1. Standard Time Series
Pre-training from scratch can be expensive, which has spurred the development of alternative approaches that leverage pre-trained models from other domains, such as large language, vision, and acoustic models. For instance, LLM4TS (Chang et al., 2023) and TEMPO (Cao et al., 2023) successfully perform time series forecasting across various datasets by fine-tuning GPT-2 (Radford et al., 2019) backbones, predicated on the notion that LLMs can be adapted to process non-linguistic datasets by activating their inherent capabilities. Similarly, Voice2Series (Yang et al., 2021) engages in the synchronization of time series and acoustic data to harness the classification prowess of an acoustic model for time series data. Another approach is presented by Wimmer et al. (Wimmer and Rekabsaz, 2023), who utilize vision-language models (VLMs) to predict market changes. Beyond fine-tuning existing models, a distinct methodology involves direct inference from LLMs for time series forecasting, showcasing commendable zero-shot performance. A notable example of this is LLMTime (Gruver et al., 2023), which introduces various strategies for effectively tokenizing time series data and transforming discrete token distributions into flexible continuous value densities.
@@
Notably, recent efforts have been directed towards creating general-purpose, single-modality standard time series foundation models. TS2Vec (Yue et al., 2022) represents a pioneering effort by introducing a universal framework for learning time series representations through contrastive learning. SimMTM (Dong et al., 2023) explores cross-domain applications, where pre-trained models via masked time series modeling exhibit superior fine-tuning performance in forecasting and classification tasks. More recent works, such as Timer (Liu et al., 2024c) and UniTS (Gao et al., 2024), further advance the field by facilitating general time series analysis through single, large-scale pre-trained models. Moreover, there is a growing interest in adapting pre-trained models, such as LLMs, for broad time series analysis applications. OFA (Zhou et al., 2023b) and TEST (Sun et al., 2023) exemplify this trend, though both approaches necessitate end-to-end fine-tuning for specific tasks.
@@ 4.2.1. Spatio-Temporal Graph
Most foundation models for spatio-temporal graphs are task-oriented and only focused on graph data. In the transportation sector, TFM (Wang et al., 2023c) utilizes graph structures and algorithms to analyze the behavior and interactions within transportation systems, showing promising results in urban traffic forecasting. ST-LLM (Liu et al., 2024b) combines spatio-temporal information with a partially frozen LLM to improve traffic predictions, while DiffSTG (Wen et al., 2023a) applies denoising diffusion models to spatio-temporal graphs for probabilistic traffic forecasting. Efforts towards domain-agnostic models include STEP (Shao et al., 2022), which links spatio-temporal GNNs with a pre-trained transformer for enhanced forecasting by learning from extensive historical data. Similarly, STGCL (Liu et al., 2022) and SPGCL (Li et al., 2022b) explore the integration of contrastive learning into spatio-temporal graph forecasting, indicating its potential benefits. Research on general-purpose models for spatio-temporal graphs is limited. A notable example, USTD (Hu et al., 2023), introduces a unified model for both forecasting and kriging tasks, employing an uncertainty-aware diffusion approach to address diverse challenges effectively.
@@ 5.1.1. Transformer-based Models
The choice of foundation model framework remains debated in the realm of time series analysis, contrasting the trend towards decoder-only models in natural language processing. Notable works in this area includes encoder-only (Nie et al., 2022; Woo et al., 2024; Gao et al., 2024), encoder-decoder (Dong et al., 2024; Garza and Mergenthaler-Canseco, 2023; Ansari et al., 2024), and decoder-only (Das et al., 2023; Rasul et al., 2023; Liu et al., 2024c) models. Ansari et al. (Ansari et al., 2024) analyze the applicability of the encoder-decoder framework to decoder-only models. Liu et al. (Liu et al., 2024c) discuss that while the encoder-only model is favored in time series forecasting for its effectiveness on small datasets, the decoder-only architecture, with its strong generalization and capacity, could be preferred for large-scale time series models. The diversity in the architectural choices underscores the potential and necessity for further exploration within this field.
@@
In terms of standard time series analysis, the Transformer architecture leverages its sequence modeling capabilities to capture temporal dynamics. This includes either repurposing pretrained LLMs for time series to leverage their preexisting sequence modeling strengths (Xue and Salim, 2022), or directly using the Transformer architecture as a base for TSFMs, training from scratch to achieve models best suited for the specifics of time series data (Garza and Mergenthaler-Canseco, 2023). Besides, various techniques have been innovated to augment the functionality of Transformer models in time series analysis comprehensively. A common practice in TSFMs segments time series into patches, which can effectively encapsulate local dynamics within input tokens (Zhou et al., 2023b; Chang et al., 2023; Das et al., 2023; Jin et al., 2023b; Rasul et al., 2023; Nie et al., 2022; Cao et al., 2023; Woo et al., 2024). Another critical design is the normalization layer, where reversible instance normalization (Kim et al., 2021) techniques, standardizing data through instance-specific mean and variance then reverting it at the output layer, have found extensive application across the above models. Moreover, specialized approaches such as multi-resolution analysis, exemplified by Moirai (Woo et al., 2024) through the employment of varying patch sizes, and decomposition strategies, as implemented by TEMPO (Cao et al., 2023) via the separation of complex interactions into the trend, seasonal, and residual components, have been shown to enhance model efficacy substantially.
@@
For spatial time series, the attention mechanism is utilized to model both the spatial and temporal dependency. For instance, ST-LLM (Liu et al., 2024b) employs a novel partially frozen attention strategy for traffic prediction, leveraging spatial-temporal embeddings to capture the intricate dynamics of traffic data across space and time. Conversely, other studies opt for independent modeling of spatial and temporal relationships. TFM (Wang et al., 2023c) is a case in point, which employs attention mechanisms within a dynamic graph encoder for spatial modeling, integrating time encoding for temporal aspects, embodying principles of transformers in addressing traffic system’s spatial-temporal dependencies. Besides simultaneously modeling spatial and temporal relationships, there exists an alternative approach that augments the Transformer model with additional spatial models or external spatial information to enhance its capabilities in the temporal modeling of time series. An example of this is STEP (Shao et al., 2022), which uses unsupervised pre-trained TransFormer blocks to model temporal relationship from long-term history time series, while applying a graph structure learner and spatio-temporal GNNs based on the representation of TransFormer blocks. Furthermore, the application of Transformer models extends to the domain of spatial-temporal prompt learning, as evidenced by initiatives such as MetePFL (Chen et al., 2023b) and FedWing (Chen et al., 2023c).
@@ 5.2.1. Pre-training
Note that the aforementioned pre-training methods typically build the model from scratch and obtain the universal knowledge from data with the same modality (i.e., time series). Nevertheless, recent advancements in time series research have heightened the usage of LLMs (Zhou et al., 2023b; Jin et al., 2023b; Xue and Salim, 2022; Chang et al., 2023; Cao et al., 2023; Gruver et al., 2023; Yu et al., 2023; Chen et al., 2023f; Xie et al., 2023; Liu et al., 2023b, 2024a, 2024b; Xue et al., 2022; Wang et al., 2023a; Shi et al., 2023; Gunjal and Durrett, 2023), VLMs (Wimmer and Rekabsaz, 2023), and AMs (Yang et al., 2021) that are pre-trained from other data modalities (text sequence, image-text sequence, acoustic signals).
@@ 5.2.2. Adaptation
Fine-tuning is a common strategy to adapt foundation models to target tasks. Based on the way the foundation model is used on the target dataset, there are three mainstream works: fine-tuning the whole model (Man et al., 2023; Nguyen et al., 2023; Pathak et al., 2022) or specific components (e.g., training positional embeddings and layer normalization, while keeping feedforward and attention layers frozen when fine-tuning LLMs) (Chang et al., 2023; Zhou et al., 2023b), to directly infer results, or integrate foundation models as part of the whole model (Sun et al., 2023; Chen et al., 2023f; Xue et al., 2022; Li et al., 2023).
@@ 5.3.1. Single-modality
A majority of current TSFMs are constructed and tailored on the basis of single-modal data. Compared with multi-modal methods, the single-modal time series modeling strategy gains the advantages of inherent simplicity and bypasses the challenges of handling modality gaps, yet frequently demonstrates excellent empirical results across a wide range of real-world applications, such as traffic (Shao et al., 2022; Liu et al., 2024b) and climate forecasting (Chen et al., 2023a; Nguyen et al., 2023).
@@ 5.3.2. Multi-modality
However, the single-modal methods may not encapsulate the full picture for several challenging downstream tasks in finance (Chen et al., 2023f; Yu et al., 2023) and healthcare domains (Liu et al., 2023b; Li et al., 2023). To cope with this issue, there have been initiatives towards developing multi-modal, task-oriented foundation models, where additional information provides useful information to enhance the model capability. In Chen et al., an external ChatGPT is queried to construct an evolving graph structure representing companies, based on the analysis of news headlines at specific time steps. As such, the inferred graph and stock prices are fed into the time series model (that uses GNN and LSTM for information propagation) to generate stock price movement predictions. Another example in healthcare also demonstrated the effectiveness of multi-modal medical context modeling, which aligns the embedding of ECG (Electrocardiogram) and corresponding medical text reports under a self-supervised contrastive learning framework and performs ECG classification. In general multi-modal time series analysis, similar cross-modality alignment strategies (e.g., contrastive learning (Sun et al., 2023), reprogramming (Jin et al., 2023b), token-wise prompting (Liu et al., 2024a)) are adopted, where the multi-modal inputs are often the textual description of datasets and pre-training word embedding from LLMs. As a notable example, Time-LLM (Jin et al., 2023b) introduces a reprogramming framework that aligns the language knowledge from pre-trained word embedding and time series information via linear projection and multi-head attention, where the handcrafted dataset descriptions are also used to quired text token embeddings as prompts, which further enhances the embedding space and informs the LLM to comprehend the task contexts. As such, utilizing multi-modal data facilitates the repurpose of the existing LLM into time series forecasters without additional computational costs.

### v2

@@ Abstract.
Q. Wen is the corresponding author. Email: qingsongedu@gmail.com
@@ 2. Background
LLM, AM, VLM [General: Time-LLM (Jin et al., 2023b), OFA (Zhou et al., 2023b), LLM4TS (Chang et al., 2023a), PromptCast (Xue and Salim, 2022),
TEMPO (Cao et al., 2023), LLMTime (Gruver et al., 2023), Voice2Series (Yang et al., 2021), AutoTimes (Liu et al., 2024b), UniTime (Liu et al., 2024a), leaf, text width=33em] [Finance: Yu et al. (Yu et al., 2023), Chen et al. (Chen et al., 2023f), Xie et al. (Xie et al., 2023), Wimmer et al. (Wimmer and Rekabsaz, 2023), leaf, text width=33em] [Healthcare: Liu et al. (Liu et al., 2023b), leaf, text width=22em] ] [Self-supervised [Generative [General: PatchTST (Nie et al., 2022), Moirai (Woo et al., 2024), Lag-Llama (Rasul et al., 2023),
TimeSiam (Dong et al., 2024), Timer (Liu et al., 2024d), Das et al. (Das et al., 2023), UniTS (Gao et al., 2024),
@@
(MLP RNN CNN) [Self-supervised [Generative [General: TSMixer (Ekambaram et al., 2023), leaf, text width=22em ] ] [Contrastive [General: TF-C (Zhang et al., [n. d.]) , TS2Vec (Yue et al., 2022) , CLUDA (Ozyurt et al., 2022) , leaf, text width=22em ] ] ] [Fully-supervised [General: TTMs (Ekambaram et al., 2024), TimesNet (Wu et al., 2022), RWKV-TS (Hou and Yu, 2024), leaf, text width=25em] ] ] [Diffusion-based [General: TimeGrad (Rasul et al., 2021) , D3VAE (Li et al., 2022a) , TransFusion (Sikder et al., 2023) , ScoreGrad (Yan et al., 2021) ,
Biloš et al. (Biloš et al., 2022) , Crabbé et al. (Crabbé et al., 2024) , TimeDiff (Shen and Kwok, 2023) , Wang et al.  (Wang et al., 2023e) , DiffTime (Coletta et al., 2024), leaf, text width=35em ] [Finance: FTS-Diffusion (Huang et al., 2024), leaf, text width=31.7em ] [Power: DiffLoad (Wang et al., 2023d), leaf, text width=31.7em ] ] ] [Spatial Time Series [Transformer-based [Pre-trained LLM [Transportation: ST-LLM (Liu et al., 2024c), TPLLM (Ren et al., 2024), leaf, text width=22em ] [General: GATGPT (Chen et al., 2023e), leaf, text width=22em ] ] [Self-supervised [Generative [Transportation: STEP (Shao et al., 2022), leaf, text width=22em] [Climate: W-MAE (Man et al., 2023), MetePFL (Chen et al., 2023b), FengWu (Chen et al., 2023a), leaf, text width=22em ] [General: UniST (Yuan et al., 2024), leaf, text width=22em] ] ] [Fully-supervised [Transportation: CPPBTR (Duan et al., 2019) , TFM (Wang et al., 2023c), leaf, text width=22em ] [Climate: FourCastNet (Pathak et al., 2022) , FedWing (Chen et al., 2023c) , Pangu-Weather (Bi et al., 2023a), ClimaX (Nguyen et al., 2023), leaf, text width=33em ] ] ] [non-Transformer-based
(MLP RNN CNN) [Self-supervised [General: SPGCL (Li et al., 2022b) , STGCL (Liu et al., 2022), leaf, text width=22em ] ] ] [Diffusion-based [General: DiffSTG (Wen et al., 2023a) , DSTPP (Yuan et al., 2023) , DYffusion (Cachay et al., 2023) , Yun et al. (Yun et al., 2023) , USTD (Hu et al., 2023) , PriSTI (Liu et al., 2023a), leaf, text width=42em ] ] ] [Others [Transformer-based [Pre-trained LLM [Mobility: AuxMobLCast (Xue et al., 2022) , LLM-Mob (Wang et al., 2023a), leaf, text width=22em ] [Event: LAMP (Shi et al., 2023) , Gunjal & Durrett et al. (Gunjal and Durrett, 2023), leaf, text width=22em ] ] [Self-supervised [Generative [Mobility: GTM (Lin et al., 2024), leaf, text width=22em] [Event: NYUTron (Jiang et al., 2023) , GatorTron (Yang et al., 2022), leaf, text width=22em ] ] [ Contrastive [Mobility: TrajCL (Chang et al., 2023b), leaf, text width=22em] ] ] ] [non-Transformer-based
@@ 4.1. Standard Time Series
Pre-training from scratch can be expensive, which has spurred the development of alternative approaches that leverage pre-trained models from other domains, such as large language, vision, and acoustic models. For instance, LLM4TS (Chang et al., 2023a) and TEMPO (Cao et al., 2023) successfully perform time series forecasting across various datasets by fine-tuning GPT-2 (Radford et al., 2019) backbones, predicated on the notion that LLMs can be adapted to process non-linguistic datasets by activating their inherent capabilities. Similarly, Voice2Series (Yang et al., 2021) engages in the synchronization of time series and acoustic data to harness the classification prowess of an acoustic model for time series data. Another approach is presented by Wimmer et al. (Wimmer and Rekabsaz, 2023), who utilize vision-language models (VLMs) to predict market changes. Beyond fine-tuning existing models, a distinct methodology involves direct inference from LLMs for time series forecasting, showcasing commendable zero-shot performance. A notable example of this is LLMTime (Gruver et al., 2023), which introduces various strategies for effectively tokenizing time series data and transforming discrete token distributions into flexible continuous value densities.
@@
Notably, recent efforts have been directed towards creating general-purpose, single-modality standard time series foundation models. TS2Vec (Yue et al., 2022) represents a pioneering effort by introducing a universal framework for learning time series representations through contrastive learning. SimMTM (Dong et al., 2023) explores cross-domain applications, where pre-trained models via masked time series modeling exhibit superior fine-tuning performance in forecasting and classification tasks. More recent works, such as Timer (Liu et al., 2024d) and UniTS (Gao et al., 2024), further advance the field by facilitating general time series analysis through single, large-scale pre-trained models. Moreover, there is a growing interest in adapting pre-trained models, such as LLMs, for broad time series analysis applications. OFA (Zhou et al., 2023b) and TEST (Sun et al., 2023) exemplify this trend, though both approaches necessitate end-to-end fine-tuning for specific tasks.
@@ 4.2.1. Spatio-Temporal Graph
Most foundation models for spatio-temporal graphs are task-oriented and only focused on graph data. In the transportation sector, TFM (Wang et al., 2023c) utilizes graph structures and algorithms to analyze the behavior and interactions within transportation systems, showing promising results in urban traffic forecasting. ST-LLM (Liu et al., 2024c) combines spatio-temporal information with a partially frozen LLM to improve traffic predictions, while DiffSTG (Wen et al., 2023a) applies denoising diffusion models to spatio-temporal graphs for probabilistic traffic forecasting. Efforts towards domain-agnostic models include STEP (Shao et al., 2022), which links spatio-temporal GNNs with a pre-trained transformer for enhanced forecasting by learning from extensive historical data. Similarly, STGCL (Liu et al., 2022) and SPGCL (Li et al., 2022b) explore the integration of contrastive learning into spatio-temporal graph forecasting, indicating its potential benefits. Research on general-purpose models for spatio-temporal graphs is limited. A notable example, USTD (Hu et al., 2023), introduces a unified model for both forecasting and kriging tasks, employing an uncertainty-aware diffusion approach to address diverse challenges effectively.
@@ 5.1.1. Transformer-based Models
The choice of foundation model framework remains debated in the realm of time series analysis, contrasting the trend towards decoder-only models in natural language processing. Notable works in this area includes encoder-only (Nie et al., 2022; Woo et al., 2024; Gao et al., 2024), encoder-decoder (Dong et al., 2024; Garza and Mergenthaler-Canseco, 2023; Ansari et al., 2024), and decoder-only (Das et al., 2023; Rasul et al., 2023; Liu et al., 2024d) models. Ansari et al. (Ansari et al., 2024) analyze the applicability of the encoder-decoder framework to decoder-only models. Liu et al. (Liu et al., 2024d) discuss that while the encoder-only model is favored in time series forecasting for its effectiveness on small datasets, the decoder-only architecture, with its strong generalization and capacity, could be preferred for large-scale time series models. The diversity in the architectural choices underscores the potential and necessity for further exploration within this field.
@@
In terms of standard time series analysis, the Transformer architecture leverages its sequence modeling capabilities to capture temporal dynamics. This includes either repurposing pretrained LLMs for time series to leverage their preexisting sequence modeling strengths (Xue and Salim, 2022), or directly using the Transformer architecture as a base for TSFMs, training from scratch to achieve models best suited for the specifics of time series data (Garza and Mergenthaler-Canseco, 2023). Besides, various techniques have been innovated to augment the functionality of Transformer models in time series analysis comprehensively. A common practice in TSFMs segments time series into patches, which can effectively encapsulate local dynamics within input tokens (Zhou et al., 2023b; Chang et al., 2023a; Das et al., 2023; Jin et al., 2023b; Rasul et al., 2023; Nie et al., 2022; Cao et al., 2023; Woo et al., 2024). Another critical design is the normalization layer, where reversible instance normalization (Kim et al., 2021) techniques, standardizing data through instance-specific mean and variance then reverting it at the output layer, have found extensive application across the above models. Moreover, specialized approaches such as multi-resolution analysis, exemplified by Moirai (Woo et al., 2024) through the employment of varying patch sizes, and decomposition strategies, as implemented by TEMPO (Cao et al., 2023) via the separation of complex interactions into the trend, seasonal, and residual components, have been shown to enhance model efficacy substantially.
@@
For spatial time series, the attention mechanism is utilized to model both the spatial and temporal dependency. For instance, ST-LLM (Liu et al., 2024c) employs a novel partially frozen attention strategy for traffic prediction, leveraging spatial-temporal embeddings to capture the intricate dynamics of traffic data across space and time. Conversely, other studies opt for independent modeling of spatial and temporal relationships. TFM (Wang et al., 2023c) is a case in point, which employs attention mechanisms within a dynamic graph encoder for spatial modeling, integrating time encoding for temporal aspects, embodying principles of transformers in addressing traffic system’s spatial-temporal dependencies. Besides simultaneously modeling spatial and temporal relationships, there exists an alternative approach that augments the Transformer model with additional spatial models or external spatial information to enhance its capabilities in the temporal modeling of time series. An example of this is STEP (Shao et al., 2022), which uses unsupervised pre-trained TransFormer blocks to model temporal relationship from long-term history time series, while applying a graph structure learner and spatio-temporal GNNs based on the representation of TransFormer blocks. Furthermore, the application of Transformer models extends to the domain of spatial-temporal prompt learning, as evidenced by initiatives such as MetePFL (Chen et al., 2023b) and FedWing (Chen et al., 2023c).
@@ 5.2.1. Pre-training
Note that the aforementioned pre-training methods typically build the model from scratch and obtain the universal knowledge from data with the same modality (i.e., time series). Nevertheless, recent advancements in time series research have heightened the usage of LLMs (Zhou et al., 2023b; Jin et al., 2023b; Xue and Salim, 2022; Chang et al., 2023a; Cao et al., 2023; Gruver et al., 2023; Yu et al., 2023; Chen et al., 2023f; Xie et al., 2023; Liu et al., 2023b, 2024b, 2024c; Xue et al., 2022; Wang et al., 2023a; Shi et al., 2023; Gunjal and Durrett, 2023), VLMs (Wimmer and Rekabsaz, 2023), and AMs (Yang et al., 2021) that are pre-trained from other data modalities (text sequence, image-text sequence, acoustic signals).
@@ 5.2.2. Adaptation
Fine-tuning is a common strategy to adapt foundation models to target tasks. Based on the way the foundation model is used on the target dataset, there are three mainstream works: fine-tuning the whole model (Man et al., 2023; Nguyen et al., 2023; Pathak et al., 2022) or specific components (e.g., training positional embeddings and layer normalization, while keeping feedforward and attention layers frozen when fine-tuning LLMs) (Chang et al., 2023a; Zhou et al., 2023b), to directly infer results, or integrate foundation models as part of the whole model (Sun et al., 2023; Chen et al., 2023f; Xue et al., 2022; Li et al., 2023).
@@ 5.3.1. Single-modality
A majority of current TSFMs are constructed and tailored on the basis of single-modal data. Compared with multi-modal methods, the single-modal time series modeling strategy gains the advantages of inherent simplicity and bypasses the challenges of handling modality gaps, yet frequently demonstrates excellent empirical results across a wide range of real-world applications, such as traffic (Shao et al., 2022; Liu et al., 2024c) and climate forecasting (Chen et al., 2023a; Nguyen et al., 2023).
@@ 5.3.2. Multi-modality
However, the single-modal methods may not encapsulate the full picture for several challenging downstream tasks in finance (Chen et al., 2023f; Yu et al., 2023) and healthcare domains (Liu et al., 2023b; Li et al., 2023). To cope with this issue, there have been initiatives towards developing multi-modal, task-oriented foundation models, where additional information provides useful information to enhance the model capability. In Chen et al., an external ChatGPT is queried to construct an evolving graph structure representing companies, based on the analysis of news headlines at specific time steps. As such, the inferred graph and stock prices are fed into the time series model (that uses GNN and LSTM for information propagation) to generate stock price movement predictions. Another example in healthcare also demonstrated the effectiveness of multi-modal medical context modeling, which aligns the embedding of ECG (Electrocardiogram) and corresponding medical text reports under a self-supervised contrastive learning framework and performs ECG classification. In general multi-modal time series analysis, similar cross-modality alignment strategies (e.g., contrastive learning (Sun et al., 2023), reprogramming (Jin et al., 2023b), token-wise prompting (Liu et al., 2024b)) are adopted, where the multi-modal inputs are often the textual description of datasets and pre-training word embedding from LLMs. As a notable example, Time-LLM (Jin et al., 2023b) introduces a reprogramming framework that aligns the language knowledge from pre-trained word embedding and time series information via linear projection and multi-head attention, where the handcrafted dataset descriptions are also used to quired text token embeddings as prompts, which further enhances the embedding space and informs the LLM to comprehend the task contexts. As such, utilizing multi-modal data facilitates the repurpose of the existing LLM into time series forecasters without additional computational costs.